Question 15 (1 point) Suppose that Spain and France both produce ships and grapes, which are sold for the same price in both countries. The table below shows the combinations of the two goods that each country carn produce in one year using the same amounts of capital and labor. Suppose that Spain is currently producing 30 ships and 30 tons of grapes, and France is currently producing 30 ships and 30 tons of grapes. If Spain and France each specialize in producing only one good (the good for which each has a comparative advantage), then how many additional ships can be produced for the two countries combined? Provide your answer as a number rounded to two decimal places. Use - to indicate negative amounts. Do not include any symbols, such as S %, or , in your answer Spain France Shi Grapes (in tons) 75 60 45 Shi Grapes (in tons) 10 15 30 45 60 75 40 30 20 10 20 40 50
Answer #1

Spain's opportunity cost for the production of ships = 75/50= 1.5 tons of grapes.

And France's opportunity cost for the production of ships = 50/75= 0.67 tons of grapes.

Because France has a lower opportunity cost in the production of ships. Therefore, France has a comparative advantage in the production of ships and Spain has a comparative advantage in the production of grapes.

If France and Spain each specialize in producing only that good in which it has a comparative advantage. Then , After trade total production of ships = 75 because France would produce it. And total production of Grapes = 75 tons because only Spain would produce it.

Before trade , the total production of ships = (30+30)= 60

And total production of grapes = (30+30)= 60

After specialization , Additional production of ships by two countries =(75-60)= 15 ships

And Additional production of grapes by two countries=(75-60)= 15 tons of grapes.